OpenStudy (anonymous):

A 1-kg model airplane has 12.5 joules of kinetic energy and 98 joules of gravitational potential energy. What is its speed? What is its height?

OpenStudy (isaidavila):

Ke=1/2(mv^2) 12.5=1/2(mv^2) 25=1*v^2 v=5

OpenStudy (isaidavila):

U=mgh 98=1*10*h 9.8=h